OpenStudy (anonymous):

Which is the equation of the line that contains points (8,10) and (-4,2)? Y-10=3/2(x-8) Y-8=3/2(x-10) Y-10=2/3(x-8) Y-8=2/3(x-10) *help I suck at math :(

OpenStudy (anonymous):

What do you think the answer is?

OpenStudy (anonymous):

you know how to find the slope?

OpenStudy (anonymous):

I have no idea how to do any of this ._.!

OpenStudy (anonymous):

(8,10) and (-4,2) \[m=\frac{y_2-y_1}{x_2-x_1}\] which in your case is \[\frac{10-2}{8-(-4)}\]as a first step

OpenStudy (anonymous):

compute this fraction get \[\frac{8}{12}\]which reduces to \(\frac{2}{3}\)

OpenStudy (anonymous):

Well we know that if a point is on the line, then it will satisfy the equation. So for your points (8,10) and (-4,2), you have two methods of solving this: 1. Sub in these values of x and y into each of the four equations and see which is the only one that works. 2. Rearrange the four equations so they are of the form y=mx+c, then calculate the line between these points and see which one of the four it matches. Take your pick! Personally, I think the first approach is easiest for this question.

OpenStudy (anonymous):

that eliminates the first two choices

OpenStudy (anonymous):

then the "point slope" formula is \[y=y_1=m(x-x_1)\] if you use the point \((8,10)\) and \(m=\frac{2}{3}\)you get \[y-10=\frac{2}{3}(x-8)\]

OpenStudy (anonymous):

Ok so I'm a try and work it out and see what I get :)

OpenStudy (anonymous):

you are not actually being asked to put in the slope intercept form, just the point slope form so you are done at that step

OpenStudy (anonymous):

Oh if see ! Thank you :)